OpenStudy (anonymous):

A software manufacturer sells one of its products at $100 per license. The costs of producing and selling the software are $9,500 for development and $5 for making a compact disc (CD), packaging, shipping, and handling. Set up the cost and revenue functions, and use this information to determine how many licenses need to be sold for the company to break even.

OpenStudy (anonymous):

cost(x) = 9500 + 5x revenue(x) = 100x breakEven is where cost(x) = revenue(x) So equate the two functions.. 100x = 9500 + 5x simplify and solve for x and you're good